# ECC Analyzer: Extract Trading Signal from Earnings Conference Calls using Large Language Model for Stock Performance Prediction

### v1

@@ 3.4 Additive Multi-modal Fusion
Given the model’s reliance on several inputs and diverse data types, we identify an effective fusion structure to integrate these features into the training process to ensure a balanced weighting among components. We use additive interactions to handle the representational fusion of different abstract representations. These operators can be viewed as differentiable building blocks that combine information from several different data streams and can be flexibly inserted into almost any unimodal pipeline Liang et al. (2022). Given the audio feature $T_{a}$, textual feature $T_{t}$ from the transcript, and $T_{s},T_{f}$ from ECC analyzed text, additive fusion can be seen as learning a new joint representation:
$E=w_{0}+w_{1}\cdot T_{a}+w_{2}\cdot T_{t}+w_{3}\cdot T_{s}+w_{4}\cdot T_{f}+\epsilon$ (5)
@@
where $w_{1}\in R^{512\times 512}$, $w_{2}\in R^{768\times 512}$ and $w_{3},w_{4}\in R^{1024\times 512}$ are the weights learned for additive fusion, $w_{0}$ the bias term and $\epsilon$ the error term. $E$ is a vector with 512 as the final feature from the Earning Conference Call Encoder.
3.5 Multi-Task Prediction
We begin our prediction process by aggregating features from various modules into a comprehensive feature representation. This unified feature set is fed into a two-layer neural network designed to perform the regression task. Integrating these diverse inputs into a cohesive output is crucial, as it harnesses the strengths of each module to enhance analysis and prediction accuracy.
Building on insights from previous research in multimodal financial risk prediction, which has demonstrated substantial improvements in prediction performance through multitask learning, we adopt a joint modeling approach. Here, we concurrently model volatility prediction and VaR prediction using a multi-task framework. The multi-task prediction module is comprised of two separate single-layer feedforward networks, each responsible for predicting volatility (vol) and Value at Risk (var) values individually. We train ECC Analyzer by optimizing multitask loss:
$\mathcal{L}$ $=\mu(\sum_{i}(\hat{y_{i}}-y_{i})^{2})$ (6)
$+(1-\mu)\text{max}(q\times(v-\hat{v}),(1-q)(\hat{v}-v))$
multi-task learning allows us to optimize performance by accurately capturing and predicting multiple stock performance factors simultaneously.
Table 1: Performance results on our proposed framework ECC Analyzer from different baseline models.
Model $\overline{MSE}$ $MSE_{3}$ $MSE_{7}$ $MSE_{15}$ $MSE_{30}$ $VaR$ Multi-Task
Classical Method 0.713 1.710 0.526 0.330 0.284 / $\otimes$
LSTM 0.746 1.970 0.459 0.320 0.235 / $\otimes$
MT-LSTM-ATT 0.739 1.983 0.435 0.304 0.233 / $\otimes$
HAN 0.598 1.426 0.461 0.308 0.198 / $\otimes$
MRDM 0.577 1.371 0.420 0.300 0.217 / $\otimes$
HTML 0.401 0.845 0.349 0.251 0.158 / D
GPT-4-Turbo 2.198 3.187 5.059 7.959 11.824 0.371 D
ECC Analyzer 0.316 0.553 0.306 0.247 0.159 0.049 D
4 Results and Discussions
4.1 Dataset
@@
The dataset utilized in this study is sourced from the publicly available S&P 500 ECC dataset as constructed by Qin and Yang (2019). It includes both audio recordings and corresponding text transcripts from the 2017 earnings calls of 500 major companies listed on the S&P 500 and traded on U.S. stock exchanges. The dataset consists of 572 unique instances where the audio recordings were accurately and closely aligned with the text transcripts. Following the setup by Qin and Yang (2019), we partitioned the dataset into a training set and a test set with an 8:2 ratio, organized temporally to ensure that the data in the training set precedes those in the test set. This temporal division is crucial for maintaining the integrity of our predictive model, aligning the training process with the principle of using historical data to predict future risks—thus enhancing the accuracy and reliability of our forecasting approach.
4.2 Baseline
We compare our approach to several important baselines including 1) GARCH-based Classical Methods; 2) LSTM Gers et al. (2000) model; 3) MT-LSTM+ATT Luong et al. (2015) employing an attention-enhanced LSTM as the foundational model; 4) HAN (Glove) uses a Hierarchical Attention Network with dual-layered attention at the word and sentence levels for prediction; 5) MRDM Qin and Yang (2019) proposed a multi-modal deep regression approach for volatility prediction tasks; 6) HTML Yang et al. (2020) presented a state-of-the-art method; and 7) GPT-4-turbo-2024-04-09: directly utilize LLM to predict stock performance. We explain each baseline method in detail in the Appendix A.1
4.3 Implementation Details
We use GPT-4 for the ECC Analyzer experiment, utilizing it to analyze ECC data and build the analysis database. Throughout this process, we set the temperature parameter to 0. This ensures that the Large Language Models (LLMs) produce the most predictable responses, which aids in maintaining consistency in our experiments.
@@ 4.3 Implementation Details
For the overall training of the ECC Analyer framework, we developed the code using PyTorch. Each Multi-Head Attention layer in the network comprises 6 layers and 8 individual heads in each layer. The training process utilized batch sizes $b\in\{2,4,8\}$. We use a grid search to determine the optimal parameters and select the learning rate $\lambda$ for Adam optimizer among $\{1e-3,1e-5,1e-6,1e-7\}$. The best hyper-parameters were kept consistent across all experiments, with the exception of the trade-off parameter $\mu$ which varied between the two tasks. We list the evaluation metrics in Appendix A.2.
4.4 Overall Results Analysis (RQ1)
Table 1 shows the performance of various methods in predicting stock performance. Notably, the ECC Analyzer framework excels, especially in short-term and medium-term forecasts, with the lowest Mean Squared Error (MSE) values. Its long-term prediction performance is comparable to the state-of-the-art method, HTML. This improvement highlights the benefits of deep analysis using LLMs extracted from ECCs. However, directly applying LLMs to stock performance prediction proves largely ineffective, similar to random guesses. This indicates that LLMs are more effective as tools to enhance investors’ understanding of a company’s financial health rather than direct predictors of financial metrics.
Additionally, the ECC Analyzer demonstrates outstanding performance in Value at Risk (VaR) prediction, confirming the efficacy of our methodology in providing a nuanced and comprehensive approach to financial risk prediction. We also report the results on Returns in Appendix B.
@@ 4.4 Overall Results Analysis (RQ1)
Table 2: Performance results of ablation study. We designed the ablation study as follows: 1) Audio+Text: uses raw audio and text data from ECCs; 2)Audio+Text+$E_{os}$: adds a comprehensive ECC summary generated by LLMs; 3) Audio+Text+$E_{os}+E_{cs}$: integrates both overall and chunk summaries for the ECC; 4) Audio+Text+$E_{fo}$: combines raw data with focused analytical results; 5) $E_{os}+E_{cs}+E_{fo}$: merges all LLM analyses for prediction without raw data; 6) Audio+Text+$E_{os}+E_{cs}+E_{fo}$: combines all data and analyses for enhanced prediction.
@@
Audio+Text+$E_{os}+E_{cs}+E_{fo}$ 0.310 0.553 0.306 0.22 0.159
4.5 Ablation Study (RQ2)
In our research, we conducted an ablation study to assess how different combinations of ECC analysis results impact our model’s performance. This systematic comparison helped us identify the individual contributions of each component.
According to Table 2, we can find that the audio and text features extracted by advanced large language models significantly improved short-term prediction accuracy compared to previous methods. Furthermore, incorporating summaries of the data slightly enhanced performance, but more notable improvements were observed when we added analysis of specific focus points. This indicates that our model effectively isolates and utilizes the most relevant information for predicting stock movements. Our best analytical results come from integrating the full spectrum of data and analytical outputs, underscoring the value of each component in our model. Notably, good predictive results were also can obtained using only analytics derived from LLMs, affirming the response to our RQ2: comprehensive analysis indeed enhances the predictive capability for stock performance.
@@ 4.5 Ablation Study (RQ2)
4.6 Interpretability Study (RQ3)
In our study, we initiate the analysis by collaborating with consultants to summarize and identify key focus points from earnings conference calls; the more comprehensive the summary, the better. Following this initial step, we combine these focus points and try different combinations to determine the optimal subset that most strongly predicts stock performance. (The finalized list of focus points is included in the appendix.) This method serves a dual purpose for investors. Firstly, it reveals which focus points are highly relevant to stock performance, allowing investors to prioritize their attention effectively. Secondly, by extracting and analyzing these key focuses, investors can understand the specific contents in the ECC that drive movements in the stock market.
5 Conclusions
This study confirms the effectiveness of using Large Language Models (LLMs) to extract and analyze key topics from earnings conference calls. Our approach not only pinpoints critical discussions but also assesses their impact on stock performance, predicting metrics like volatility and Value at Risk (VaR). By integrating textual and audio data, our model offers a comprehensive view, capturing subtleties such as tone and pitch. Results from benchmark comparisons demonstrate our model’s superior accuracy and predictive capabilities, highlighting LLMs’ potential to improve interpretability and decision-making in investments.
Appendix A Supplement Material for Experiment Setup
A.1 Baseline Setup
  • Classical Methods: We incorporate the GARCH model and its derivatives, as described in  Franses and Van Dijk (1996); Kim and Won (2018). These models are well-recognized for short-term volatility prediction but may not be as effective for forecasting average volatility over longer periods, such as n-day volatility.
  • LSTM Gers et al. (2000): Long Short-Term Memory Networks (LSTMs) are a popular choice for financial time series prediction due to their efficacy in handling sequential data. We use a straightforward LSTM model as a benchmark for volatility prediction.
  • MT-LSTM+ATT Luong et al. (2015): combines the prediction of average n-day volatility with the forecasting of single-day volatility, employing an attention-enhanced LSTM as the foundational model.
  • HAN (Glove): uses a Hierarchical Attention Network with dual-layered attention at the word and sentence levels. HAN first get word embeddings using pre-trained Glove vectors and then processed by a Bi-GRU Chung et al. (2014) encoder, while another Bi-GRU encoder simultaneously forms a sentence-level representation of each document. The resulting document representation is input into a regression layer to produce predictions.
  • MRDM Qin and Yang (2019): The MRDM model first introduced a multi-modal deep regression approach to fuse the GloVe embeddings and hand-crafted acoustic features for volatility prediction tasks.
  • HTML Yang et al. (2020): This work presented a state-of-the-art model that employs WWM-BERT for text token encoding. Similar to MDRM, HTML also leverages the same audio features. These unimodal features are then combined and processed through a sentence-level transformer, resulting in multimodal representations for each call.
  • GPT-4-turbo-2024-04-09: We assessed the capability of LLMs in directly predicting stock performance from ECCs. The model was set to generate response with a zero temperature setting to ensure deterministic output.
A.2 Evalation Metrics
A.3 Volatility of Different Future Intervals
We used the following formula to calculate the Volatility of Different Terms:
$\sigma_{ijk}^{2}=\frac{1}{k-1}\sum_{m=1}^{k}[S_{i,(j-m)}-\frac{1}{k}\sum_{l=1}^{k}S_{i,(j-l)}]^{2}$
Here, we define the $\sigma_{ijk}$ as the $k-days$ volatility of Stock $i$ at time $j$, which is calculated as the sample standard deviation of the past $k-days$ stock closed price of company $i$. But in the real application, we took the log form of historical volatility.
A.4 VaR of Different Future Intervals
Our second task is predicting the 1-day VaR of the target stock based on the multi-source inputs. The definition of VaR is:
$x=F^{-1}(p)$
The $F(\cdot)$ is the cumulative loss distribution, $p$ is the percentile we set, and $x$ is the VaR. From the idea of Quantile Regression, we can have:
$L_{\tau}(y,\hat{y})=\begin{cases}\tau\cdot(y-\hat{y})&\text{if }y\geq\hat{y}\\ (1-\tau)\cdot(\hat{y}-y)&\text{if }y<\hat{y}\end{cases}$
Calculating and estimating VaR can help the company better deal with financial risks and avoid extreme scenarios in the future.
A.5 Returns with Different Future Intervals
We used the following formula to calculate the Returns with Different Gaps:
$R_{ijk}=\frac{S_{i,(j+k)}-S_{i,j}}{S_{i,j}}$
Here, we denote $S_{i,j}$ as the closed price of company $i$’s stock at time $j$, and $R_{ijK}$ as the $k-day$ return of stock price $i$ at time $j$. We calculated this value as the difference between the closed price of company $i$ at time $j$ and time $j+k$ divided by the closed price of company $i$ at time $j$.
Appendix B The Results on Returns
To evaluate the predictive capability of our model, we have applied it to forecast stock returns. The results of these forecasts are systematically compiled and displayed in a table 3.
Table 3: ECC Analyzer’s Prediction Performance On Stock Returns
Return_3d Return_7d Return_15d Return_30d
Predited Return 0.0024 0.0068 0.0131 0.0185
True Return 0.0007 0.0017 0.0035 0.0139
Error 0.0018 0.0051 0.0096 0.0046
Percentage Error 2.6923 3.0964 2.7216 0.3348
Appendix C Regression and Sentiment Analysis
In the main chapter, we used the various returns and volatility of different terms as our predicted target. This part will introduce how we calculated these values and further results about the relationships between sentiments calculated by LLM and these target values.
C.1 Cumulative Abnormal Returns and Regression Analysis
We used both Abnormal Returns collected from the market and the returns of different terms to capture the sentiment delivered by ECCs. For the abnormal Returns, we used the following regression model Lee and Connolly (2010):
$R_{it}=\alpha+\beta R_{mt}+\varepsilon_{it}$
$R_{it}$ is the return of the security at time $t$, also $R_{mt}$ is the market portfolio return at time $t$. The Abnormal Returns we defined here are the difference between the real value of Returns and estimated ones, which is $\varepsilon_{it}$. For other dependent variables, we used the $R_{ijk}$ for $k=1,3,7,15,30$. Results are shown in Table 4.
Table 4: This table shows the linear regression results of different kinds of returns. We used different types of returns as dependent variables. $CARs$ is the cumulative abnormal returns, calculated as the difference between the actual value of the predicted valve from factor models. The rest of the dependent variables are calculated as the difference between the closed price of company $i$ at time $j$ and time $j+k$ divided by the closed price of company $i$ at time $j$. The independent variable is the Sentiment score, which is calculated by LLM. We extracted all the events in one ECC and evaluated their sentiment score. The percentage of events with positive sentiment minus the percentage of events with negative sentiments calculates the score.
Dependent Variable
CARs Return_1d Return_3d Return_7d Return_15d Return_30d
Constant $-0.0105^{***}$ $-0.0105^{***}$ $-0.0126^{***}$ $-0.0182^{***}$ $-0.0286^{***}$ $-0.0244^{***}$
(0.004) (0.004) (0.004) (0.005) (0.007 (0.008)
Sentiment $0.0223^{***}$ $0.024^{***}$ $0.0295^{***}$ $0.0442^{***}$ $0.0716^{***}$ $0.0852^{***}$
(0.008) (0.008) (0.009) (0.010) (0.015) (0.016)
Adjusted $R^{2}$ 0.016 0.018 0.022 0.038 0.048 0.055
Observation 572 572 572 572 572 572
For the volatility, we used the difference between the two closed periods of volatility. The dependent variable for volatility could be calculated as $log(\sigma_{i(j+1)k})-log(\sigma_{ijk})$ for $k=3,7,15,30$. Results are shown in Table 5.
Table 5: This table shows the linear regression results of different terms of Volatility. We used four different terms of volatility. Volatility here is defined by the log form of the sample standard deviation of the past $k-days$ stock closed price of company $i$. The independent variable is the Sentiment score, which is calculated by LLM. We extracted all the events in one ECC and evaluated their sentiment score. The percentage of events with positive sentiment minus the percentage of events with negative sentiments calculates the score.
Dependent Variable
Volatility_3d Volatility_7d Volatility_15d Volatility_30d
Constant $-0.0254^{***}$ $-0.017^{***}$ $-0.0071^{***}$ $-0.005^{***}$
(0.013) (0.009) (0.006) (0.004)
Sentiment $-0.0681^{***}$ $-0.0556^{***}$ $-0.0438^{***}$ $-0.0266^{***}$
(0.026) (0.018) (0.013) (0.009)
Adjusted $R^{2}$ 0.013 0.019 0.022 0.019
Observation 572 572 572 572
From Table 4 and Table 5, we can find that the Sentiment generated from the LLM is always statistically significant in the linear regression model results. Comparing the coefficient of different terms of Returns, we found that the coefficient gets larger when there are more significant term gaps. This indicates that the impact of the ECC’s sentiment will appear as time progresses. There is a higher probability that the impact of the ECCs occurs after a certain period because, besides the financial data, most information, like future projects mentioned in ECCs, will not be reflated instantly. That could explain why the coefficient will rise as the period gets larger. However, the volatility change will have a smaller impact when the period gets larger(because the absolute value of Sentiments’ coefficient in Table 5 is smaller when the period gets more extensive), which means that the same ECCs sentiments effect will have a more significant impact in short terms rather than the longer terms.
Figure 3: We asked four experts to help us analyze the possible scenarios of ECCs. Our experts focus on the following general Categories of the ECCs’ transcript: Financial Indicators, Employee & Managers, Cost, Project Expansion, Business, and Future Look. Each general category is separated into some minor focusing points. Our experts put forward three questions for each small focusing point. We used these questions to ask LLM and wished for a directional response in the thinking pattern hidden in the questions that we wanted to teach the LLM.
Appendix D Prompt Design
D.1 Prompt for Summarizing Earnings Conference Call Segments

### v2

@@ 4.4. Multimodal Fusion and Prediction
Given the model’s reliance on several inputs and diverse data types, we identify an effective fusion structure to integrate these features into the training process to ensure a balanced weighting among components. We use additive interactions to handle the representational fusion of different abstract representations. These operators can be viewed as differentiable building blocks that combine information from several different data streams and can be flexibly inserted into almost any unimodal pipeline (Liang et al., 2022). Given the audio feature $T_{a}$, textual feature $T_{t}$ from the transcript, and $T_{s},T_{f}$ from ECC analyzed text, additive fusion can be seen as learning a new joint representation:
(9) $E=w_{0}+w_{1}\cdot T_{a}+w_{2}\cdot T_{t}+w_{3}\cdot T_{s}+w_{4}\cdot T_{f}+\epsilon$
@@
where $w_{1}\in R^{512\times 512}$, $w_{2}\in R^{768\times 512}$ and $w_{3},w_{4}\in R^{1024\times 512}$ are the weights learned for additive fusion, $w_{0}$ the bias term and $\epsilon$ the error term. $E$ is a vector with 512 as the final feature. This unified feature set $E$ is fed into two fully connected layers to perform the regression task. We train ECC Analyzer by optimizing the MSE loss:
(10) $\mathcal{L}$ $=\mu(\sum_{i}(\hat{y_{i}}-y_{i})^{2})$
5. Experiment
We describe the datasets used in the experiment and the baseline settings for performance comparisons. Our experiments aim to answer the following research questions (RQs): RQ1: Can our proposed pipeline, when integrated with LLMs, significantly improve volatility prediction performance compared to current state-of-the-art (SOTA) approaches? RQ2: Does the synergy between our model and the LLM result in better performance compared to raw LLM predictions? RQ3: Does each distinct component derived from ECC contribute to improving volatility prediction accuracy?
5.1. Dataset
@@
The dataset utilized in this study is sourced from the publicly available S&P 500 ECC dataset as constructed by (Qin and Yang, 2019). It includes both audio recordings and corresponding text transcripts from the 2017 earnings calls of 500 major companies listed on the S&P 500 and traded on U.S. stock exchanges. The dataset consists of 572 unique instances where the audio recordings were accurately and closely aligned with the text transcripts. Following the setup by (Qin and Yang, 2019), we partitioned the dataset into a training set and a test set with an 8:2 ratio, organized temporally to ensure that the data in the training set precedes those in the test set. This temporal division is crucial for maintaining the integrity of our predictive model, aligning the training process with the principle of using historical data to predict future risks—thus enhancing the accuracy and reliability of our forecasting approach.
5.2. Baseline Setup
We compare our approach to several important baselines including:
  • Classical Method: We incorporate the GARCH model and its derivatives, as described in  (Franses and Van Dijk, 1996; Kim and Won, 2018). This model is well-recognized for short-term volatility prediction but may not be as effective for forecasting average volatility over longer periods, such as n-day volatility.
  • LSTM (Gers et al., 2000): The Long Short-Term Memory (LSTM) based method is a popular choice for financial time series prediction due to its efficacy in handling sequential data. We use a straightforward LSTM model as a benchmark for volatility prediction.
  • MT-LSTM-ATT (Luong et al., 2015): combines the prediction of average n-day volatility with the forecasting of single-day volatility, employing an attention-enhanced LSTM as the foundational model.
  • HAN (Glove): uses a Hierarchical Attention Network with dual-layered attention at the word and sentence levels. HAN first gets word embeddings using pre-trained Glove vectors and then processed by a Bi-GRU (Chung et al., 2014) encoder, while another Bi-GRU encoder simultaneously forms a sentence-level representation of each document. The resulting document representation is input into a regression layer to produce predictions.
  • MRDM (Qin and Yang, 2019): The MRDM model first introduced a multi-modal deep regression approach to fuse the GloVe embeddings and hand-crafted acoustic features for volatility prediction tasks.
  • HTML (Yang et al., 2020): This work presented a state-of-the-art model that employs WWM-BERT for text token encoding. Similar to MDRM, HTML also leverages the same audio features. These unimodal features are then combined and processed through a sentence-level transformer, resulting in multimodal representations for each call.
  • AMA-LSTM (Wang et al., 2024): The paper uses a strategy of adversarial training to minimize the effect of speaker gender in earnings conference calls.
  • GPT-4-turbo-2024-04-09: We assessed the capability of LLMs in directly predicting volatility performance from ECCs. The model was set to generate a response with a zero temperature setting to ensure deterministic output.
5.3. Implementation Details
In the experiment, all interactions with LLMs were conducted using “GPT-4 Turbo-2024-04-09”. We set the temperature parameter to 0 to ensure that the LLMs produce the most predictable responses, thereby maintaining consistency in our experiments.
@@ 5.3. Implementation Details
For the overall training of the ECC Analyer framework, we developed the code using PyTorch. Each Multi-Head Self-Attention layer in the model comprises 6 layers and 8 individual heads in each layer. The training process utilized batch sizes $b\in\{2,4,8,16\}$. We use a grid search to determine the optimal parameters and select the learning rate $\lambda$ for Adam optimizer among $\{1e-3,1e-5,1e-6,1e-7\}$.
5.4. Overall Results Analysis (RQ1 & RQ2)
Table 1. Performance results on our proposed framework ECC Analyzer with different baseline methods.
Model $\overline{MSE}$ $MSE_{3}$ $MSE_{7}$ $MSE_{15}$ $MSE_{30}$
Classical Method 0.713 1.710 0.526 0.330 0.284
LSTM 0.746 1.970 0.459 0.320 0.235
MT-LSTM-ATT 0.739 1.983 0.435 0.304 0.233
HAN (Glove) 0.598 1.426 0.461 0.308 0.198
MRDM 0.577 1.371 0.420 0.300 0.217
HTML 0.401 0.845 0.349 0.251 0.158
AMA-LSTM / 0.680 0.360 0.230 /
GPT-4-Turbo 2.198 3.187 5.059 7.959 11.824
ECC Analyzer 0.314 0.553 0.306 0.237 0.158
@@ 5.4. Overall Results Analysis (RQ1 & RQ2)
Table 2. Performance results of ablation study. We designed the ablation study as follows: 1) Audio+Text: uses raw audio and text data from ECCs; 2)Audio+Text+$E_{os}$: adds an overall ECC summary generated by LLMs; 3)Audio+Text+$E_{cs}$: adds the chucks summary generated by LLMs; 4) Audio+Text+$E_{os}+E_{cs}$: integrates both overall and chunk summaries for the ECC; 5) Audio+Text+$E_{fo}$: combines raw data with focused analytical results; 6) $E_{os}+E_{cs}+E_{fo}$: merges all LLM analyses for prediction without raw data; 6) Audio+Text+$E_{os}+E_{cs}+E_{fo}$: combines all data and analyses for enhanced prediction.
@@
Audio+Text+$E_{os}+E_{cs}+E_{fo}$ 0.314 0.553 0.306 0.237 0.158
Table 1 shows the performance of various methods in predicting volatility performance. Notably, the ECC Analyzer framework excels, especially in short-term forecasts (day 3 and day 7), with the lowest Mean Squared Error (MSE) values of 0.553 and 0.306, respectively. Its long-term prediction performance is comparable to the state-of-the-art method, HTML. However, the ECC Analyzer performs slightly lower than AMA-LSTM in predicting medium-term (day 15) volatility. Nevertheless, ECC Analyzer’s overall average MSE achieves superior performance. These encouraging experimental results illustrate that extracting fine-grained information from ECC data using LLMs can significantly enhance the model’s volatility prediction accuracy. In particular, the extracted fine-grained information features are especially beneficial for improving short-term forecasting performance, which is crucial for investors. In summary, addressing RQ1, the prediction performance of our proposed method surpasses that of the current SOTA approaches.
In response to RQ2, directly applying LLMs to volatility prediction proves largely ineffective, akin to random guessing. These results also raise concerns about the potential misuse of LLMs, considering their impact on public safety. If investors use LLMs inappropriately for numerical outputs, they may increase financial risk. This indicates that LLMs are more effective as tools to enhance investors’ understanding of a company’s financial health rather than direct predictors of financial metrics.
5.5. Ablation Study (RQ3)
In our research, we conducted an ablation study to assess how different combinations of ECC analysis results impact our model’s performance. This systematic comparison helped us identify the individual contributions of each component. Each component is represented as follows: $E_{os}$ represents the embedding of the overall ECC summary; $E_{cs}$ represents the embedding of the summary for chunks; $E_{cs}$ represents the embedding of the extracted fine-grained important sentence.
According to Table 2, we can find that the audio and text features extracted by advanced large language models significantly improved short-term prediction accuracy compared to previous methods. Furthermore, incorporating summaries of the data slightly enhanced performance, but more notable improvements were observed when we added analysis of specific focus points. This indicates that our model effectively isolates and utilizes the most relevant information for predicting stock movements. Our best analytical results come from integrating the full spectrum of data and analytical outputs, underscoring the value of each component in our model. We also obtain good predictive results using only analytics derived from LLMs, affirming the response to our RQ3: comprehensive analysis indeed enhances the predictive capability for stock volatility prediction performance.
@@ 5.5. Ablation Study (RQ3)
6. Conclusion
In this paper, we propose the ECC Analyzer, a novel framework that leverages large language models (LLMs) for in-depth analysis of ECC data to enhance volatility prediction performance. The ECC Analyzer extracts information from ECCs at multiple levels to assist in volatility prediction. Our experiments demonstrate that our proposed method improves the overall average Mean Squared Error (MSE). Specifically, we achieve better results in short-term forecasting, with medium-term and long-term forecasts also on par with SOTA methods. Additionally, we analyze the role of each component in the proposed framework for the forecasting process through ablation experiments, highlighting the contribution of each element to the overall performance.
Appendix A Appendix
A.1. Prompt Design
A.1.1. Prompt for Summarizing Earnings Conference Call Segments